OpenStudy (anonymous):

Need a little help with this formula. y=mx+b. x1=2, y1=-3, m=-1/5. I have so far y=-1/5x?? what is the rest of it?

OpenStudy (anonymous):

y+3=(-1/5)(x-2)

OpenStudy (anonymous):

Expand y=mx+b to (y-k)=m(x-h) where (h,k) is a point. From the given information, you know x=2 and y=-3, so (h,k) is (2, -3). Check signs when plugging in.

OpenStudy (anonymous):

Afterwards, distribute your (-1/5) into (x-2), subtract 3 on both sides, and viola. :)

OpenStudy (anonymous):

I got the y+3=-1/5(x-2). So subtract 3 form both sides is y=-1/5x+1?

OpenStudy (anonymous):

Watch the parenthesis. You're going to get y=(-1/5)x+2/5-3. y=-x/5-13.5 is your final equation.

OpenStudy (anonymous):

I thought that the -1/5x would not change. I am lost now.

OpenStudy (anonymous):

I gave that as an answer but it says it is wrong.

OpenStudy (anonymous):

ok, so here's what happened. (-1/5)x means (-1/5) times x, which is the same as putting x over a -5. (-1/5)x=-x/5. I'd review a little over notation, if I were you. :)

OpenStudy (anonymous):

y=-x/5-13/5, I meant.

OpenStudy (anonymous):

Ok got it. that worked thank you.